# Local multiplicity for fractional linear equations with Hardy potentials

### v1

@@ 1.2. Main results
satisfying (G) and satisfying the following property: equation (1.2) admits a nontrivial radial weak solution $u\in\mathcal{D}^{s,2}_{\textup{rad}}(\mathbb{R}^{N})$ such that
@@
There exists a radial potential $\mathscr{V}\in C^{1}(B_{1}\setminus\{0\})\cap C^{0}(B_{1})$ satisfying (G) and satisfying the following property: there is a nontrivial radial weak solution $u\in\mathcal{D}^{s,2}_{\textup{rad}}(\mathbb{R}^{N})$ to
@@ 4. Non-uniqueness
thus the proof concludes by suitably scaling. Indeed, by choosing $u(x):=u_{i}(\bar{r}x)$, $\mathscr{V}(x):=\bar{r}^{2s}\mathscr{V}_{i}(\bar{r}x)$ and $R:=\bar{r}^{-1}$, we have that $u$ satisfies (1.2) and (1.5) and that $\mathscr{V}$ belongs to $C^{1,\alpha}_{\textup{loc}}(B_{1}\setminus\{0\})\cap L^{q}(B_{1})$ and satisfies (G). Notice that $u\in H^{s}(\mathbb{R}^{N})$, $u\equiv 0$ outside $B_{R}$ since $u_{i}\equiv 0$ outside $B_{1}$. Moreover, if $\Omega$ is a nonempty open set in $B_{R}$ then $u$ cannot vanish identically in $\Omega$: indeed, if $u\equiv 0$ in an open set $\Omega\subseteq B_{R}$, then $u_{i}\equiv 0$ in an open subset of $B_{1}$. At this point, since $u_{i}$ weakly solves (1.6), by Theorem 4.1 we would conclude that $u_{i}\equiv 0$ in $\mathbb{R}^{N}$, and this is a contradiction.
@@
where $\mathscr{V}_{0}:=h(u_{1},u_{2,\delta})$. We now claim that $\mathscr{V}_{0}\in C^{1,\alpha}_{\textup{loc}}(B_{\rho}\setminus\{0\})\cap L^{q}(B_{\rho})$ ( dovrebbe andare bene con lo stesso $\alpha$ ), for some $\rho\in(0,R_{\delta}]$, and that $\mathscr{V}_{0}$ satisfies (G). More precisely, $\rho\in(0,R_{\delta}]$ is chosen in such a way that $u_{1},u_{2,\delta}>0$ in $B_{\rho}\setminus\{0\}$ (the existence of such $\rho$ is ensured by (4.4)). As a first step, we observe that, since $g\in C^{\infty}(\mathbb{R}\setminus\{0\})$, then $h\in C^{\infty}((0,+\infty)\times(0,+\infty))$ and, as a consequence, since $u_{1},u_{2,\delta}\in C^{1,\alpha}_{\textup{loc}}(B_{\rho}\setminus\{0\})$ and $u_{1},u_{2,\delta}>0$ in $B_{\rho}\setminus\{0\}$, then $\mathscr{V}_{0}\in C^{1,\alpha}_{\textup{loc}}(B_{\rho}\setminus\{0\})$ as well. We now prove that $\mathscr{V}_{0}\in L^{q}(B_{\rho})$. In view of Lagrange theorem, for any $t_{1},t_{2}>0$ there exists $t_{12}>0$ such that
@@
In this way, since $\mathscr{V}_{0}\in C^{1,\alpha}_{\textup{loc}}(B_{\rho}\setminus\{0\})\cap L^{q}(B_{\rho})$ and since $\mathscr{V}_{0}$ satisfies (G), we get $\mathscr{V}\in C^{1,\alpha}_{\textup{loc}}(B_{1}\setminus\{0\})\cap L^{q}(B_{1}),$ $\mathscr{V}$ satisfies (G) as well, and $u$ satisfies (1.5) and (1.2). Moreover, $u\equiv 0$ outside $B_{R}$ since $w\equiv 0$ outside $B_{1}$, and having shown that $w$ cannot vanish identically on a nonempty open subset of $B_{1}$, we obtain that $u$ cannot vanish identically on a nonempty open subset of $B_{R}$.
@@
This proof closely follows the one of Theorem 1.1, hence we only sketch it. In the case $\lambda=0$, we have $\gamma_{1}^{s}(0)=0$, where $\gamma_{1}^{s}(0)$ is given by (3.8), see also Lemma 3.6. Let $0<\alpha<\tfrac{4s}{N-2s}$. If $N\geq 6s$, let $p\in(\alpha+2,2^{*}_{s})$ while, if $N<6s$, let $p\in(3,2^{*}_{s})$. As in the proof of Theorem 1.1, let $u_{1},u_{2}\in H^{s}_{0,\textup{rad}}(B_{1})$ be two distinct, nontrivial solutions of (1.6), whose existence is ensured by Proposition 2.1. In view of Theorem 3.3 we have that, for $i=1,2$, $u_{i}\in C^{1}(B_{1})$ and there exist $\gamma_{1},\gamma_{2}\in[0,+\infty)$ such that (4.1) holds for some $\kappa_{i}\neq 0$ (it is not restrictive to assume $\kappa_{i}>0$), so that there exist $c_{i}\in\{0,\kappa_{i}\}$ such that $u_{i}(0)=c_{i}$, with $c_{i}=0$ if and only if $\gamma_{i}>0$. Radiality implies $u^{\prime}_{i}(0)=0$.
Let us consider the case in which $u_{i}(0)=0$ (for either $i=1$ or $i=2$). By (4.1) we can find $\bar{r}\in(0,1]$ such that $u_{i}>0$ in $B_{\bar{r}}\setminus\{0\}$. Letting $\mathscr{V}_{i}:=|u_{i}|^{p-2}$, one can easily check that $\mathscr{V}_{i}\in C^{1}(B_{\bar{r}}\setminus\{0\})$ , that $\mathscr{V}_{i}\in C^{0,\alpha}_{\textup{loc}}(B_{1})$ if $N\geq 6s$ (since $p-2>\alpha$) , that $\mathscr{V}_{i}\in C^{1}(B_{1})$ if $N<6s$ and that $\mathscr{V}_{i}$ satisfies (G). One concludes by reasoning as in the proof of Theorem 1.1, choosing $u(x):=u_{i}(\bar{r}x)$, $\mathscr{V}(x)=\bar{r}^{2s}\mathscr{V}_{i}(\bar{r}x)$ and $R:=\bar{r}^{-1}$. In this way, one has indeed $\mathscr{V}\in C^{1}(B_{1}\setminus\{0\})$, $\mathscr{V}\in C^{0,\alpha}_{\textup{loc}}(B_{1})$ if $N\geq 6s$, $\mathscr{V}\in C^{1}(B_{1})$ if $N<6s$, $u(0)=0$, $(-\Delta)^{s}u=\mathscr{V}u$ in $B_{1}$, $u\equiv 0$ outside $B_{R}$ and $u\not\equiv 0$ on any open set contained in $B_{R}$
@@
so that (4.4) holds with $\gamma=0$, and finally let $w:=u_{1}-u_{2,\delta}$. Moreover, let $R_{\delta}:=\delta^{-\frac{p-2}{2s}}\leq 1$ and $\mathscr{V}_{0}:=h(u_{1},u_{2,\delta}),$ where $h$ is defined in (4.6). Since $h\in C^{0,p-2}_{\textup{loc}}(\mathbb{R}^{2})$ if $p\leq 3$ (corresponding to the case $N\geq 6s$) and $h\in C^{1}(\mathbb{R}^{2})$ if $p>3$ (corresponding to the case $N<6s$), and since $u_{1},u_{2,\delta}\in C^{1}(B_{R_{\delta}})$, we notice that $\mathscr{V}_{0}\in C^{0,\alpha}_{\textup{loc}}(B_{R_{\delta}})$ if $N\geq 6s$ and $\mathscr{V}_{0}\in C^{1}(B_{R_{\delta}})$ if $N<6s$. We also have $\mathscr{V}_{0}\in C^{1}(B_{\rho})$ for some suitable $\rho\in(0,R_{\delta}]$, even if $N\geq 6s$, since a small enough $\rho$ can be chosen such that $u_{1},u_{2,\delta}>0$ in $B_{\rho}$, in view of (4.4). Moreover, by following the very same argument as in the proof of Theorem 1.1, one can check that $\mathscr{V}_{0}$ satisfies the condition (G). By construction, we have that $w\in H^{s}_{0,\textup{rad}}(B_{1})$ satisfies
@@
Acknowledgments
The authors warmly thank Lorenzo Brasco and Veronica Felli for useful discussions and suggestions. E.M. is supported by the MIUR- PRIN project 202244A7YL. R.O. is supported by the European Research Council (ERC), through the European Union’s Horizon 2020 project ERC VAREG - Variational approach to the regularity of the free boundaries (grant agreement No. 853404) and by the 2024 INdAM-GNAMPA project no. CUP_E53C23001670001. R.O. also acknowledges the MIUR Excellence Department Project awarded to the Department of Mathematics, University of Pisa, CUP I57G22000700001. B.V. is supported by the MIUR-PRIN project 2022SLTHCE. This manuscript reflects only the authors’ views and opinions and the Ministry cannot be considered responsible for them. The authors are members of the GNAMPA group of the Istituto Nazionale di Alta Matematica.

### v2

@@ 1.2. Main results
satisfying (G), such that $\mathscr{V}>0$ in $B_{1}\setminus\{0\}$ and satisfying the following property: equation (1.2) admits a nontrivial radial weak solution $u\in\mathcal{D}^{s,2}_{\textup{rad}}(\mathbb{R}^{N})$ such that
@@
There exists a radial potential $\mathscr{V}\in C^{1}(B_{1}\setminus\{0\})\cap C^{0}(B_{1})$ satisfying (G), such that $\mathscr{V}>0$ in $B_{1}\setminus\{0\}$ and satisfying the following property: there is a nontrivial radial weak solution $u\in\mathcal{D}^{s,2}_{\textup{rad}}(\mathbb{R}^{N})$ to
@@ 4. Non-uniqueness
thus the proof concludes by suitably scaling. Indeed, by choosing $u(x):=u_{i}(\bar{r}x)$, $\mathscr{V}(x):=\bar{r}^{2s}\mathscr{V}_{i}(\bar{r}x)$ and $R:=\bar{r}^{-1}$, we have that $u$ satisfies (1.2) and (1.5) and that $\mathscr{V}$ belongs to $C^{1,\alpha}_{\textup{loc}}(B_{1}\setminus\{0\})\cap L^{q}(B_{1})$, satisfies (G) and $\mathscr{V}>0$ in $B_{1}\setminus\{0\}$. Notice that $u\in H^{s}(\mathbb{R}^{N})$, $u\equiv 0$ outside $B_{R}$ since $u_{i}\equiv 0$ outside $B_{1}$. Moreover, if $\Omega$ is a nonempty open set in $B_{R}$ then $u$ cannot vanish identically in $\Omega$: indeed, if $u\equiv 0$ in an open set $\Omega\subseteq B_{R}$, then $u_{i}\equiv 0$ in an open subset of $B_{1}$. At this point, since $u_{i}$ weakly solves (1.6), by Theorem 4.1 we would conclude that $u_{i}\equiv 0$ in $\mathbb{R}^{N}$, and this is a contradiction.
@@
where $\mathscr{V}_{0}:=h(u_{1},u_{2,\delta})$. We now claim that, for some $\rho\in(0,R_{\delta}]$, $\mathscr{V}_{0}\in C^{1,\alpha}_{\textup{loc}}(B_{\rho}\setminus\{0\})\cap L^{q}(B_{\rho})$, $\mathscr{V}>0$ in $B_{\rho}\setminus\{0\}$ and that $\mathscr{V}_{0}$ satisfies (G). More precisely, $\rho\in(0,R_{\delta}]$ is chosen in such a way that $u_{1},u_{2,\delta}>0$ in $B_{\rho}\setminus\{0\}$ (the existence of such $\rho$ is ensured by (4.4)). As a first step, we observe that, since $g\in C^{\infty}(\mathbb{R}\setminus\{0\})$, then $h\in C^{\infty}((0,+\infty)\times(0,+\infty))$ and, as a consequence, since $u_{1},u_{2,\delta}\in C^{1,\alpha}_{\textup{loc}}(B_{\rho}\setminus\{0\})$ and $u_{1},u_{2,\delta}>0$ in $B_{\rho}\setminus\{0\}$, then $\mathscr{V}_{0}\in C^{1,\alpha}_{\textup{loc}}(B_{\rho}\setminus\{0\})$ as well. Moreover, since $g$ is increasing, we have $\mathscr{V}_{0}>0$ in $B_{\rho}\setminus\{0\}$. We now prove that $\mathscr{V}_{0}\in L^{q}(B_{\rho})$. In view of Lagrange theorem, for any $t_{1},t_{2}>0$ there exists $t_{12}>0$ such that
@@
In this way, since $\mathscr{V}_{0}\in C^{1,\alpha}_{\textup{loc}}(B_{\rho}\setminus\{0\})\cap L^{q}(B_{\rho})$, $\mathscr{V}_{0}>0$ in $B_{\rho}\setminus\{0\}$ and since $\mathscr{V}_{0}$ satisfies (G), we get $\mathscr{V}\in C^{1,\alpha}_{\textup{loc}}(B_{1}\setminus\{0\})\cap L^{q}(B_{1})$, $\mathscr{V}>0$ in $B_{1}\setminus\{0\}$, $\mathscr{V}$ satisfies (G) as well, and $u$ satisfies (1.5) and (1.2). Moreover, $u\equiv 0$ outside $B_{R}$ since $w\equiv 0$ outside $B_{1}$, and having shown that $w$ cannot vanish identically on a nonempty open subset of $B_{1}$, we obtain that $u$ cannot vanish identically on a nonempty open subset of $B_{R}$.
@@
This proof closely follows the one of Theorem 1.1, hence we only sketch it. In the case $\lambda=0$, we have $\gamma_{1}^{s}(0)=0$, where $\gamma_{1}^{s}(0)$ is given by (3.8), see also Lemma 3.6. Let $0<\alpha<\tfrac{4s}{N-2s}$. If $N\geq 6s$, let $p\in(\alpha+2,2^{*}_{s})$ while, if $N<6s$, let $p\in(3,2^{*}_{s})$. As in the proof of Theorem 1.1, let $u_{1},u_{2}\in H^{s}_{0,\textup{rad}}(B_{1})$ be two distinct, nontrivial solutions of (1.6), whose existence is ensured by Proposition 2.1. In view of Theorem 3.3 and Corollary 3.5 we have that, for $i=1,2$, $u_{i}\in C^{1}(B_{1})$ and there exist $\gamma_{1},\gamma_{2}\in[0,+\infty)$ such that (4.1) holds for some $\kappa_{i}\neq 0$ (it is not restrictive to assume $\kappa_{i}>0$), so that there exist $c_{i}\in\{0,\kappa_{i}\}$ such that $u_{i}(0)=c_{i}$, with $c_{i}=0$ if and only if $\gamma_{i}>0$. Radiality implies $u^{\prime}_{i}(0)=0$.
Let us consider the case in which $u_{i}(0)=0$ (for either $i=1$ or $i=2$). By (4.1) we can find $\bar{r}\in(0,1]$ such that $u_{i}>0$ in $B_{\bar{r}}\setminus\{0\}$. Letting $\mathscr{V}_{i}:=|u_{i}|^{p-2}$, one can easily check that $\mathscr{V}_{i}\in C^{1}(B_{\bar{r}}\setminus\{0\})$, $\mathscr{V}_{i}>0$ in $B_{\bar{r}}\setminus\{0\}$, that $\mathscr{V}_{i}\in C^{0,\alpha}_{\textup{loc}}(B_{1})$ if $N\geq 6s$ (since $p-2>\alpha$), that $\mathscr{V}_{i}\in C^{1}(B_{1})$ if $N<6s$ and that $\mathscr{V}_{i}$ satisfies (G). One concludes by reasoning as in the proof of Theorem 1.1, choosing $u(x):=u_{i}(\bar{r}x)$, $\mathscr{V}(x)=\bar{r}^{2s}\mathscr{V}_{i}(\bar{r}x)$ and $R:=\bar{r}^{-1}$. In this way, one has indeed $\mathscr{V}\in C^{1}(B_{1}\setminus\{0\})$, $\mathscr{V}>0$ in $B_{1}\setminus\{0\}$, $\mathscr{V}\in C^{0,\alpha}_{\textup{loc}}(B_{1})$ if $N\geq 6s$, $\mathscr{V}\in C^{1}(B_{1})$ if $N<6s$, $u(0)=0$, $(-\Delta)^{s}u=\mathscr{V}u$ in $B_{1}$, $u\equiv 0$ outside $B_{R}$ and $u\not\equiv 0$ on any open set contained in $B_{R}$
@@
so that (4.4) holds with $\gamma=0$, and finally let $w:=u_{1}-u_{2,\delta}$. Moreover, let $R_{\delta}:=\delta^{-\frac{p-2}{2s}}\leq 1$ and $\mathscr{V}_{0}:=h(u_{1},u_{2,\delta}),$ where $h$ is defined in (4.6). Since $h\in C^{0,p-2}_{\textup{loc}}(\mathbb{R}^{2})$ if $p\leq 3$ (corresponding to the case $N\geq 6s$) and $h\in C^{1}(\mathbb{R}^{2})$ if $p>3$ (corresponding to the case $N<6s$), and since $u_{1},u_{2,\delta}\in C^{1}(B_{R_{\delta}})$, we notice that $\mathscr{V}_{0}\in C^{0,\alpha}_{\textup{loc}}(B_{R_{\delta}})$ if $N\geq 6s$ and $\mathscr{V}_{0}\in C^{1}(B_{R_{\delta}})$ if $N<6s$. We also have $\mathscr{V}_{0}\in C^{1}(B_{\rho})$ for some suitable $\rho\in(0,R_{\delta}]$, even if $N\geq 6s$, since a small enough $\rho$ can be chosen such that $u_{1},u_{2,\delta}>0$ in $B_{\rho}$, in view of (4.4). In addition, by monotonicity of $g$ we get that $\mathscr{V}_{0}>0$ in $B_{\rho}\setminus\{0\}$. Moreover, by following the very same argument as in the proof of Theorem 1.1, one can check that $\mathscr{V}_{0}$ satisfies the condition (G). By construction, we have that $w\in H^{s}_{0,\textup{rad}}(B_{1})$ satisfies
@@
Acknowledgments. The authors warmly thank Lorenzo Brasco and Veronica Felli for useful discussions and suggestions. E.M. is supported by the MIUR- PRIN project 202244A7YL. R.O. is supported by the European Research Council (ERC), through the European Union’s Horizon 2020 project ERC VAREG - Variational approach to the regularity of the free boundaries (grant agreement No. 853404) and by the 2024 INdAM-GNAMPA project no. CUP_E53C23001670001. R.O. also acknowledges the MIUR Excellence Department Project awarded to the Department of Mathematics, University of Pisa, CUP I57G22000700001. B.V. is supported by the MIUR-PRIN project 2022SLTHCE. This manuscript reflects only the authors’ views and opinions and the Ministry cannot be considered responsible for them. The authors are members of the GNAMPA group of the Istituto Nazionale di Alta Matematica.
Data availability. Data sharing not applicable to this article as no datasets were generated or analyzed during the current study.
Declarations
Conflict of interest. On behalf of all authors, the corresponding author states that there is no Conflict of interest